OpenStudy (mrs.ambrose614):

@will.h (04.05) The price of gold has increased by 35% per year from 2000. In the year 2000, Harry bought a gold ring for $590. Which of the following functions f(x) can be used to represent the price of the ring x years after 2000? f(x) = 590(1.35)x f(x) = 590(0.65)x f(x) = 35(0.41)x f(x) = 35(1.59)x

OpenStudy (jerry45):

i cant help sry!

OpenStudy (will.h):

exponential growth form f(x) = p(1 + r)^x where x is the time and r is rate and p is principal f(x) = 590(1 + 0.35)^x f(x) = 590(1.35)^x

OpenStudy (mrs.ambrose614):

So the awnser is a

OpenStudy (will.h):

yeah